OpenStudy (anonymous):

. What is the remainder of(x^2+2x+14)÷(x+7) A. 49 B. 7 C. 2 D. -5

OpenStudy (anonymous):

(x^2+2x+14)÷(x+7) Let x+7=0 i.e. x=-7 Let us put x=-7 in the given expression we find (-7)^2+2(-7)+14 =49-14+14=49 Thus 49 is the required remainder Hence option A 49 is the correct answer. @LaBlueEyes2013